In this part, calculate the interest rates. Use the Excel Rate function to compute the interest rates.

  1. In 2018, one of the first copper pennies struck at the Philadelphia mint in 1793 was sold for $300,000. What was the rate of return on this investment? (6 Points)

  1. Calculate the interest rates in the table below using the Excel Rate function. (6 Points)

Present Value

Years

Future Value

Interest Rate

$750

10

$1,380

?

$900

12

$1,750

?

$16,000

25

$142,800

?

$70,500

16

$312,900

?
Answer #1

B А 1 Part A 2 PV 3 FV 4 NPER 5 Rate 6 7 8 Part B Present $0.01 $300,000 225 7.95% 9 Value Years 10 11 12 13 $750 $900 $16,001 Part A 2 PV 3 FV 4 NPER 5 Rate 0.01 300000 =2018-1793 =RATE(B4,,-B2,B3) 8 Part B 9 Present Value Years Future Value Interes